# CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?

### v1

@@ 5 Future Work
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All datasets, task templates, and evaluation procedures are documented in the main text and appendix. An anonymous repository containing the full source code, experiment scripts, and detailed reproduction instructions has been made publicly available at: https://github.com/Peiyu-Georgia-Li/crochetBench. This ensures that all reported results can be independently verified and extended by the research community.
Appendix A Related Work
A.1 Multi-modal Datasets Beyond Captioning
Most multimodal benchmarks have centered on descriptive pairing of images and natural language. Large-scale resources such as COCO [lin2015microsoftcococommonobjects] and Flickr30k [plummer2016flickr30kentitiescollectingregiontophrase] provide dense captions of everyday scenes, advancing vision–language representation learning. More recent datasets extend beyond captioning to procedural or instructional domains. Recipe1M+ aligns food images with ingredient lists and cooking steps [marin2019recipe1mdatasetlearningcrossmodal], while instructional video corpora such as YouCook2 [zhou2017automaticlearningproceduresweb] and HowTo100M [miech2019howto100mlearningtextvideoembedding] pair narrated demonstrations with visual segments. These resources emphasize semantic alignment but generally evaluate with retrieval- or similarity-based metrics.
Our benchmark departs from this paradigm by pairing images with executable procedures. Rather than asking models to generate a semantically similar description, we require them to synthesize a program (CrochetPARADE DSL) that can be rendered and structurally verified. This shift enables functional evaluation—akin to program synthesis—and reduces reliance on subjective similarity measures.
A.2 Bridging Visual Reasoning and Procedural Language
Procedural understanding benchmarks highlight the importance of sequential, state-dependent reasoning. Datasets such as Assembly101 [sener2022assembly101largescalemultiviewvideo] and IKEA-ASM [benshabat2023ikeaasmdatasetunderstanding] capture human assembly activities, modeling dependencies across actions, objects, and preconditions. Instructional video benchmarks further test long-horizon understanding and error detection [wei2025instructionbenchinstructionalvideounderstanding, lin2024videoguibenchmarkguiautomation]. Our task complements this line of work by grounding supervision not in temporally segmented actions but in artifact-centric procedures—finished crochet items paired with symbolic, stepwise instructions. This enables models to reason about topology, geometry, and sequential dependencies in a single unified representation.
A.3 Crafting and Domain-Specific Procedural Data
Closer to our domain, prior work has begun to explore fiber crafts. Seitz et al. introduced Digital Crochet, a visual, graph-based notation system for representing crochet patterns [digital_crochet]. In knitting, Kaspar et al. developed Neural Inverse Knitting, mapping images of knitted patterns to machine instructions [kaspar2019neuralinverseknittingimages]. These works underscore the feasibility of executable supervision in crafts, but remain limited in scale and scope. Our benchmark builds on this foundation by providing thousands of real, community-tested crochet patterns, along with a compiler and renderer for executable evaluation.
A.4 Relation to Code Benchmarks
The executable nature of CrochetPARADE connects it directly to program synthesis benchmarks. HumanEval [chen2021evaluatinglargelanguagemodels], MBPP [austin2021programsynthesislargelanguage], and APPS [hendrycks2021measuringcodingchallengecompetence] evaluate code generation by execution against unit tests. Similarly, Spider [yu2019spiderlargescalehumanlabeleddataset] and CoNaLa [yin2018mining] frame natural language to code translation tasks. In vision, Im2LaTeX-100K [deng2017imagetomarkupgenerationcoarsetofineattention] and pix2code [beltramelli2017pix2codegeneratingcodegraphical] evaluate image-to-program translation with render fidelity as the metric. Our tasks—ranging from text-to-DSL translation to image-to-stitch recognition—extend this paradigm to the crafting domain, introducing structured 2D/3D topology as part of the evaluation. This grounds multimodal learning in a setting where success requires both semantic alignment and structural correctness.
A.5 Domain-Specific Languages for Executable Evaluation
A key enabler of our benchmark is the use of a domain-specific language. Prior DSL-based benchmarks, such as Im2LaTeX and the CROHME competition on handwritten math recognition [crohme], demonstrate how symbolic formalisms enable reproducible rendering and structural evaluation. CrochetPARADE adopts this principle for fiber crafts: each pattern compiles into an abstract syntax tree that can be rendered and tested, supporting metrics such as ExecPass@K, structural unit tests, and image–render similarity. This functional perspective moves beyond surface-level similarity to test whether a model’s output actually works, in contrast to chemistry and recipe domain, which require quite amount of time to validate it [huang2024application, chefFusion, huang2025chemhgnn].
@@ B.1 Instruction Complexity by Skill Level
Example Dataset Entry
Table 10: Representative pattern entry from CrochetBench.
@@
Instructions (truncated for brevity)
B.2 Skill Level Distribution
B.2.1 Overall Distribution
Table 11 summarizes the overall distribution of skill levels across the CrochetBench dataset. The majority of patterns are labeled as easy (58.7%), followed by intermediate (32.3%). Only a small fraction are classified as beginner (7.6%) or experienced (1.3%). [^3] Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern). Together they account for $0.1\%$ of the dataset.
@@ B.2.1 Overall Distribution
B.2.2 Distribution by Project Type
@@
B.3 Pattern Complexity Analysis
B.3.1 Instruction Length Statistics
@@
B.3.2 Abbreviation Statistics
@@
B.3.3 Complexity by Skill Level
@@
Table 15: Instruction length and abbreviation counts by skill level.
@@
B.3.4 Most and Least Complex Project Types
Finally, we identify the most complex and simplest project types by average instruction length. Tables 16 and 17 list the top 10 categories. Garments such as dresses, vests, pants, and tunics are the most demanding, with average instructions exceeding 5,800 characters. By contrast, smaller accessories such as cowls, washcloths, scarves, and headbands are substantially shorter, typically under 2,000 characters.
@@ C.5 Task D (Project-Level): NL $\rightarrow$ DSL Translation Prompt
NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Convert instructions + images into compilable CrochetPARADE DSL code. Output only the DSL code. No explanations, commentary, or extra text. Example 1: ”image path”: https://www.yarnspirations.com/cdn/shop/files/BRC0116-035467M.jpg, INSTRUCTIONS Note: Join with sl st to first sc at end of each rnd. Ch 2. **Rnd 1:** 6 sc in 2nd ch from hook. Join. (6 sc) **Rnd 2:** Ch 1. 2 sc in each sc around. Join. (12 sc) **Rnd 3:** Ch 1. (2 sc in next sc, 1 sc in next sc) repeat around. End with 1 sc. Join. (18 sc) **Rnd 4:** Ch 1. (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (24 sc) **Rnd 5:** Ch 1. Sc in each sc around. Join. (24 sc) **Rnd 6:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (30 sc) **Rnds 7–8:** Repeat Rnd 5 (sc in each sc). Join. (30 sc each round) **Rnd 9:** Ch 1. **Working in back loops only**: (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (40 sc) **Rnd 10:** Ch 1. Sc in each sc around (both loops). Join. (40 sc) **Rnd 11:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (50 sc) **Finish:** Fasten off. DSL ¶ch.B ¶sc@B.A,5sc@B,ss@A ¶ch.A,sk,6sc2inc,ss@A ¶ch.A,sk,[sc2inc,sc]*6,ss@A ¶ch.A,sk,[sc2inc,2sc]*6,ss@A ¶ch.A,sk,24sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*6,ss@A ¶[ch.A,sk,30sc,ss@A ¶]*2 ¶ch.A,sk,[scbl,scbl@[@],2scbl]*10,ss@A ¶ch.A,sk,40sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*10,ss@A USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
@@ Appendix D CrochetPARADE: Pattern Renderer, Analyzer, and Debugger
CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D [crochetparade]. By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions. [^4] https://www.crochetparade.org/
@@ Appendix E DSL Error Taxonomy
To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples. Errors fall into four major groups:
1. Syntax Structure Errors
@@
Multiple References Without Parentheses.
@@
2. Stitch Definition Errors
Stitch Not Found.
@@ 2. Stitch Definition Errors
3. Labeling and Reference Errors
@@
4. Structural and Formatting Issues
Turning Errors.
@@ 4. Structural and Formatting Issues
Variable Naming Issues.
Conflicts between variable names and stitch names.

### v2

@@ 6 Conclusion
Appendix A Crochet Primer
Crochet patterns describe how to construct a textile artifact through a sequence of symbolic stitch instructions. Each instruction specifies an operation performed with a hook and yarn, and the resulting pattern is defined by the order, repetition, and spatial arrangement of these stitches. This appendix summarizes only the conventions needed to interpret the examples in our benchmark.
Basic stitch types.
Crochet relies on a small vocabulary of atomic stitches, each producing a loop with a characteristic height and structure. The most common stitches in U.S. notation are:
  • ch (chain): foundational stitch used to begin rows or rounds.
  • sc (single crochet): a short, dense stitch.
  • hdc (half double crochet) and dc (double crochet): taller stitches that build height more quickly.
  • sl st (slip stitch): a joining stitch used for connecting motifs or closing rounds.
  • bobble (bobble stitch): a cluster of 5 partially completed double crochet stitches closed together into a single stitch.
These stitches can be combined in rows (worked back and forth) or rounds (worked in a circle).
Pattern syntax and structure.
Crochet instructions follow a compact symbolic notation. A pattern is organized into rows or rounds, each specifying a sequence of stitches. For example:
Row 3: Ch 1, sc in each st across, turn.
Instructions may include:
  • Repetition: indicated by parentheses and a multiplier, e.g., (sc, ch 1) 3 times.
  • Increases/decreases: e.g., 2 sc in next st (increase) or sc2tog (single-crochet two stitches together; decrease).
  • Stitch counts: patterns often end rows or rounds with “—N sc,” indicating the number of stitches that should remain.
Relationship to symbolic representations.
Each crochet instruction corresponds to a local modification of the fabric’s topology. This makes crochet patterns naturally suited to symbolic or program-like representations such as CrochetPARADE, which encode stitches as structured primitives with explicit control flow (loops, groups, labels). Because stitch sequences fully determine the geometry of the final artifact, correctness can be assessed by verifying the structure of the generated program or by rendering the corresponding stitch graph.
This primer covers the minimal terminology required to interpret our dataset and evaluation tasks. For readers interested in additional background, standard crochet references provide extended stitch catalogs and diagram conventions.
@@ Appendix B Additional Dataset Statistics
CrochetBench supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments. Figure 6 lists the ten most common categories by frequency. The majority of patterns belong to a small number of dominant types—Afghans and Blankets alone account for over one-quarter of the dataset.
Each pattern is labeled with one of four primary skill levels, including beginner, easy, intermediate, or experienced. This allows for stratified evaluation across complexity tiers. Figure 7 shows the skill level distribution, which is strongly skewed toward beginner-friendly content. More details can be found at Appendix B. Instructional complexity varies substantially across patterns. The number of characters in each instruction ranges from 20 to over 30,000, with a mean of 3,216 and a median of 2,453. Abbreviation counts (i.e., unique stitch tokens per pattern) range from 1 to 31, with an average of 10.6. These statistics are summarized in Appendix B.1.
We observe a clear correlation between skill level and instruction length: beginner patterns tend to be short and use fewer abbreviations, while experienced patterns are significantly longer and more symbolically dense.
In addition to symbolic complexity, the dataset contains 3,143 abbreviation instances mapped to 789 unique standardized stitch tokens. This lexical mapping enables tasks such as vocabulary translation, sequence generation, and instruction validation. Beyond raw instructions, the structured schema also records rich metadata. A representative dataset entry is provided in Table 4.
@@ B.1 Instruction Complexity by Skill Level
Table 3 shows the instruction complexity by skill level.
@@
Figure 7: Skill level distribution across the CrochetBench dataset.
B.2 Example Dataset Entry
Table 4 demonstrate a crochet dataset example.
@@ B.2 Example Dataset Entry
Instructions Instructions: Note: Ch 2 at beg of each rnd counts as hdc. BAG With MC, ch 4. Join with sl st to form ring. 1st rnd: Ch 2. 11 hdc in ring. Join with sl st to top of ch 2. 12 hdc. 2nd rnd: Ch 2. 1 hdc in same sp as sl st. 2 hdc in each hdc around. Join. 24 hdc. 3rd rnd: Ch 2. 1 hdc in same sp. 1 hdc in next hdc. *2 hdc in next hdc, 1 hdc in next.* Rep around. Join. 36 hdc. 4th rnd: Ch 2. 1 hdc in each hdc around. Join. 5th rnd: Ch 2. 1 hdc in same sp. 1 hdc in next 2 hdc. *2 hdc, 1 hdc in next 2.* Join. 48 hdc. 6th rnd: As 4th rnd. 7th rnd: Ch 2. 1 hdc in next 2 hdc. *2 hdc, 1 hdc in next 3.* Rep. 60 hdc. 8th rnd: Ch 2. Back loops only, 1 hdc around. Join. 9th–13th rnds: Ch 2. 1 hdc in each hdc around. Join. 14th rnd: Ch 2. 1 hdc in same sp. 1 hdc in next 4 hdc. *2 hdc, 1 hdc in next 4.* Join. 72 hdc. 15th rnd: Ch 2. 1 hdc in same sp. 1 hdc in next 5 hdc. *2 hdc, 1 hdc in next 5.* Join. 84 hdc. 16th–22nd rnds: Ch 2. 1 hdc in each hdc around. Join. 23rd rnd: Ch 2. 1 hdc in next 4 hdc. *Hdc2tog, 1 hdc in next 5.* Rep. Hdc2tog. Join. 72 sts. 24th rnd: Ch 2. 1 hdc in next 3 hdc. *Hdc2tog, 1 hdc in next 4.* Rep. Join. 60 sts. 25th rnd: Ch 2. 1 hdc in next 2 hdc. *Hdc2tog, 1 hdc in next 3.* Rep. Join. 48 sts. Fasten off. Eyes (Make 2) With A, ch 8. 1st rnd: 2 sc in 2nd ch from hook. 1 sc in next 5 ch. 3 sc in last ch. Continue on rem loops, 1 sc in each ch. Join. 17 sc. 2nd rnd: Ch 1. 3 sc in first sc. 1 sc in next 7 sc. 3 sc in next sc. 1 sc in next 8 sc. Join. Fasten off. Handle With MC, ch 45. 1st row: 1 sc in 2nd ch from hook. 1 sc across. 44 sc. Turn. 2nd row: Ch 1. 1 sc across. Turn. Rep last row 4 more times. Fasten off. Finishing Sew Eyes to Bag. Embroider mouth and teeth with A. Attach Handle. Cut cardboard circle to fit bottom.
Table 4: Representative pattern entry from CrochetBench.
B.3 Skill Level Distribution
B.3.1 Overall Distribution
Table 5 summarizes the overall distribution of skill levels across the CrochetBench dataset. The majority of patterns are labeled as easy (58.7%), followed by intermediate (32.3%). Only a small fraction are classified as beginner (7.6%) or experienced (1.3%).***Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern). Together they account for $0.1\%$ of the dataset.
@@ B.3.1 Overall Distribution
B.3.2 Distribution by Project Type
@@
B.4 Pattern Complexity Analysis
B.4.1 Instruction Length Statistics
@@
Table 7: Instruction length statistics (in characters).
B.4.2 Abbreviation Statistics
@@
Table 8: Abbreviation count statistics.
B.4.3 Complexity by Skill Level
@@
Table 9: Instruction length and abbreviation counts by skill level.
B.4.4 Most and Least Complex Project Types
Finally, we identify the most complex and simplest project types by average instruction length. Tables 10 and 11 list the top 10 categories. Garments such as dresses, vests, pants, and tunics are the most demanding, with average instructions exceeding 5,800 characters. By contrast, smaller accessories such as cowls, washcloths, scarves, and headbands are substantially shorter, typically under 2,000 characters.
@@ C.5 Task D (Project-Level): NL $\rightarrow$ DSL Translation Prompt
NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Convert instructions + images into compilable CrochetPARADE DSL code. Output only the DSL code. No explanations, commentary, or extra text. Example 1: ”image path”: https://www.yarnspirations.com/cdn/shop/files/BRC0116-035467M.jpg, INSTRUCTIONS Note: Join with sl st to first sc at end of each rnd. Ch 2. **Rnd 1:** 6 sc in 2nd ch from hook. Join. (6 sc) **Rnd 2:** Ch 1. 2 sc in each sc around. Join. (12 sc) **Rnd 3:** Ch 1. (2 sc in next sc, 1 sc in next sc) repeat around. End with 1 sc. Join. (18 sc) **Rnd 4:** Ch 1. (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (24 sc) **Rnd 5:** Ch 1. Sc in each sc around. Join. (24 sc) **Rnd 6:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (30 sc) **Rnds 7–8:** Repeat Rnd 5 (sc in each sc). Join. (30 sc each round) **Rnd 9:** Ch 1. **Working in back loops only**: (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (40 sc) **Rnd 10:** Ch 1. Sc in each sc around (both loops). Join. (40 sc) **Rnd 11:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (50 sc) **Finish:** Fasten off. DSL ¶ch.B ¶\$k=0\$,sc@B.A[k],5sc@B,ss@A[k++] ¶ch.A[k],sk,6sc2inc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,sc]*6,ss@A[k++] ¶ch.A[k],sk,[sc2inc,2sc]*6,ss@A[k++] ¶ch.A[k],sk,24sc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,3sc]*6,ss@A[k++] ¶[ch.A[k],sk,30sc,ss@A[k++] ¶]*2 ¶ch.A[k],sk,[scbl,scbl@[@],2scbl]*10,ss@A[k++] ¶ch.A[k],sk,40sc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,3sc]*10,ss@A[k++] ¶ch.A[k],sk,[2sc,3sc]*10,ss@A[k++] USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
@@ Appendix D CrochetPARADE: Pattern Renderer, Analyzer, and Debugger
CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D (Tassev, 2025). By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions.***https://www.crochetparade.org/
@@ Appendix E DSL Error Taxonomy
To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples.
@@
Multiple References.
@@
Undefined Stitch.
@@
Variable Naming Conflict.
Conflicts between variable names and stitch names.
@@
Turning Issue.